The population of scores on the SAT is normally distributed with a µ = 500 and σ = 100. If you were to take a random sample of 48 students who had taken the SAT, what are the chances that their mean would be less than 520?
Answer #1

Solu tiau population scores.し then PL S20]520-500 oo/48 1o/u 0,57 ,4162. 162
Answer #2

SOLUTION :


SD of sample = SD of population / sqrt( sample size)

= 100/sqrt(48) 

= 14.4338 


P( x < 520) = P(z < (520 - 500)/14.4338) = P(z < 1.3856)


= 0.9162 + (0.9177 - 0.9162)* 56/100

= 0.9170 

= 91.70 % 


So, chances to be less than 520 score = 91.70 % (ANSWER).


answered by: Tulsiram Garg